OpenStudy (anonymous):

Help 2x + 1 greater than or equal to 5 or -3x > -9

OpenStudy (jhannybean):

\[\large 2x+1 \ge 5 \ \ \quad \text{or} \ \ -3x > -9\]like this?

OpenStudy (anonymous):

Yes

OpenStudy (jhannybean):

Solve it part by part. \[\large 2x+1 \ge 5\] subtract -1 from both sides. \[\large 2x \ge 4\] divide by 2.\[\large x \ge 2\] now the next one. \[\large -3x > -9\] divide both sides by -3. Remember, when dividing by a negative, flip your sign. \[\large \frac{-3x}{-3}<\frac{-9}{-3} \]\[\large x < 3\] We can put them together such as : \(\large 2 \le x <3\)

OpenStudy (rmrjr22):

bam!

OpenStudy (jhannybean):

Understand? :)

OpenStudy (anonymous):

Yes Thank you (: